OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 2 points) Evaluate (5 + 2)2 − 40 + 3 6 11 12 34

OpenStudy (nubeer):

first solve the parts in bracket. then apply D.M.A.S or Bodmas.. it's simple arithmetic.

OpenStudy (theviper):

\((5 + 2)^2 − 40 + 3\) =49-40+3=12 12 is the answer :)

Parth (parthkohli):

\[(5 + 2)^2 - 40 + 3\]First, solve for the parentheses. 5 + 2 = 7. Now the exponent. What's 7^2?

OpenStudy (nubeer):

@ParthKohli we were talking about something a while ago and lol it happened again..